OpenStudy (anonymous):

How do I find the slope in this equation -2x - 4y = 10

OpenStudy (anonymous):

In y = mx + b form, the slope is in the m position. So lets put your equation in y = mx + b form... -2x - 4y = 10 --- add 2x to both sides -4y = 2x + 10 --- divide both sides by -4 y = 2/-4x + 10/-4 y = -1/2x - 5/2 Now what number is in the m position ? That number is your slope.